OpenStudy (anonymous):

Find the discriminat of 8x^2+5x+6=0

OpenStudy (anonymous):

217

OpenStudy (anonymous):

sorry -167

OpenStudy (campbell_st):

for aa quadratic in the form \[ax^2 + bx + c = 0\] the dicriminant is \[b^2 - 4 \times a \times c\] in your question a = 8, b = 5 and c = 6 substitute and calculate...

OpenStudy (anonymous):

the equation has no real roots